# Complementary polynomials in quantum signal processing

### v1

@@ Abstract
Quantum signal processing is a framework for implementing polynomial functions on quantum computers. To implement a given polynomial $P$, one must first construct a corresponding complementary polynomial $Q$. Existing approaches to this problem employ numerical methods that are not amenable to explicit error analysis. We present a new approach to complementary polynomials using complex analysis. Our main mathematical result is a contour integral representation for a canonical complementary polynomial. The integral representation on the unit circle has a particularly simple and efficacious Fourier analytic interpretation, which we use to develop a Fast Fourier Transform-based algorithm for the efficient calculation of $Q$ in the monomial basis with explicit error guarantees. Numerical evidence that our algorithm outperforms the state-of-the-art optimization-based method for computing complementary polynomials is provided.
@@ 1 Introduction
Quantum signal processing (QSP) [1] and its extensions [2, 3] describe a simple single-qubit parameterised circuit that applies a chosen polynomial to a scalar. They have become indispensable in state-of-the-art quantum algorithms because the circuits can be lifted from scalars to arbitrary matrices, called quantum singular value transformation (QSVT) [4, 5]. The matrix is embedded inside a larger unitary, called a block encoding. The polynomial is then applied to all singular values of the matrix simultaneously, which can lead to a quantum speedup compared to classical evaluation of matrix functions. QSVT has had tremendous impact in quantum computing since polynomials may be used to approximate a wide variety of functions. Thereby, this family of algorithms encompasses many prior quantum algorithms [6], including those for Hamiltonian simulation [7], solving linear systems [8], phase estimation [9], and amplitude amplification [10], often even improving the prior algorithm.
There are different parameterizations of QSP [1, 4, 11]. The fundamental idea underlying each is that certain polynomials may be realized within a matrix element of a finite product of single-qubit unitaries [1, 3, 12]. Determining the parameters of the unitaries required to implement a particular polynomial has proven to be a challenging problem for which several numerical methods have been developed [11, 13, 14, 15, 16]. The generalized quantum signal processing (GQSP) framework [^1] For future convenience, in particular to relate (1.2) to other QSP conventions, we have written (1.2) as the transpose of [3, Eq. 7]. of Motlagh and Wiebe [3] has, as we describe below, clarified the mathematical structure of this factorization problem.
@@
In (1.2), we call $Q$ a complementary polynomial to $P$ and the parameters $\lambda$, $(\phi_{j})_{j=0}^{d}$, $(\theta_{j})_{j=0}^{d}$ phase factors. It is shown in [3] that once a complementary polynomial is known, it is mathematically trivial to construct the phase factors to realize (1.2). This remarkable fact stands in contrast to the paradigm of that for standard QSP where, with or without knowledge of the complementary polynomial, one must recourse to numerical methods to determine the phase factors. GQSP subsumes the Laurent [11] and real formulations of QSP, as we show in Appendix A; see also [7]. Thus, the conceptual advantage of GQSP, that the challenge in realizing any polynomial lies wholly in computing a complementary polynomial, descends to standard QSP. Further, the phase factors required in lifted algorithms operating on matrices like QSVT follow immediately from those in QSP. By the above discussion, the fundamental obstacle to representing polynomials in both the GQSP and standard QSP as well as QSVT frameworks is the following problem.
@@
Previous approaches to this problem rely on root-finding [4], Prony’s method [15], or optimization [3]. In contrast, we construct the first exact representation of a canonical complementary polynomial $Q$, valid throughout the entire complex plane, in the form of a set of contour integrals. The contour integral representation of $Q$ on $\mathbb{T}$ can be rephrased in the language of Fourier analysis. We combine this Fourier analytic interpretation and the fast Fourier transform to develop efficient numerical algorithms to compute $Q$ in the monomial basis.
Exact and explicit error analysis of our algorithms is performed, providing a rigorous upper bound on the classical runtime as a linear function, up to logarithmic factors, of the error and degree. Numerical results from our reference implementation further demonstrate the practicality and competitiveness of our algorithm. We emphasize that existing numerical approaches to the construction of complementary polynomials, which we describe in Section 1.3 below, rely on heuristics and so are not amenable to such error analysis.
In the remainder of this introduction, we state our principal results, give remarks on their proofs, describe related literature, introduce notation used in the main text, and outline the structure of the paper.
@@ 1.1 Statement of results
the restriction that $p_{0}\neq 0$ is imposed without loss of generality as $\lvert z^{n}P(z)\rvert=\lvert P(z)\rvert$ holds for all $z\in\mathbb{T}$, $n\in\mathbb{Z}$. Our principal result is the following theorem, giving representations of $Q$ on $\mathbb{D}\coloneqq\{z\in\mathbb{C}:\lvert z\rvert<1\}$, $\mathbb{T}$, and $\mathbb{C}\setminus\overline{\mathbb{D}}$.
@@
the integration contour $\mathbb{T}$ is positively-oriented, and the dashed integral indicates a principal value prescription with respect to the singularity $z^{\prime}=z$ on $\mathbb{T}$, solves Problem 1. Moreover, (1.4c) is, up to a multiplicative phase, the unique solution of Problem 1 with no roots in $\mathbb{D}$.
Theorem 2 provides an exact representation of $Q$ in Problem 1, in canonical form: all roots lie outside of $\mathbb{D}$. In the important special case where $P$ has real coefficients, a canonical complementary polynomial with real coefficients exists. The precise statement, a mild generalization of a result in [13, Section 8], is now given.
@@
Numerical methods
Corollary 2.2 suggests a practical numerical method to compute an approximation of the complementary polynomial, supposing $Q_{0}$ is known. This is trivially the case if a $\delta>0$ is known with $||P(z)||_{\infty,\mathbb{T}}<1-\delta$; then $Q_{0}(z)=1$. In other cases, $Q_{0}(z)=1$ can be attained by first scaling $P(z)$ down slightly, $P(z)\to(1-\varepsilon)P(z)$ with a small $\varepsilon>0$.
First, for cases with a bound $||P(z)||_{\infty,\mathbb{T}}\leq(1-\delta)$, Algorithm 1 solves Problem 1 in time $O(N\log N)$ using the fast Fourier transform (FFT), where the even parameter $N\in\mathbb{Z}_{\geq d}$ defines the discrete Fourier basis $\{\mathrm{e}^{\mathrm{i}n\theta}\}_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$. An informal description of our algorithm based on Corollary 2.2 with $Q_{0}(z)=1$ is as follows:
@@
    Compute approximations to the Fourier coefficients $(a_{n})_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$ of $\log\big{(}1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}\big{)}$ using the FFT, in time $O(N\log N)$.
@@
    Compute approximations to $Q$ at the $N$th roots of unity by applying the Fourier multiplier $\Pi$ in Fourier space, (1.6)–(1.8), using an FFT and inverse FFT in time $O(N\log N)$.
@@
    Compute approximations to the coefficients of $Q/{Q}_{0}$ in the monomial basis using the result of the previous step and the FFT, in time $O(N\log N)$.
We prove in Theorem 3 that this algorithm is efficient, with a sufficient $N$ asymptotically scaling as $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$.
In all other cases, when $||P(z)||_{\infty,\mathbb{T}}=1$ or $\delta$ is so small that the runtime $N\sim 1/\delta$ proven for Algorithm 1 is undesirable, Algorithm 2 solves Problem 1. It first scales down the input polynomial slightly before running Algorithm 1 in time $O(N\log N)$. Theorem 4 proves that a sufficient $N$ has asymptotic scaling $O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$, eradicating any dependence on $\delta$.
We give a reference implementation of our algorithm and provide numerical evidence that it outperforms the optimization-based approach to complementary polynomials from [3] in Section 5.
@@ 1.2 Remarks on the results and their proofs
The following remarks apply to Theorem 2 and Corollary 2.2.
@@
    The crucial observation leading to Theorem 2 is that the real part of the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately, can be determined exactly on $\mathbb{T}$ using (1.1). This allows for the use of the Schwarz integral formula [17] to determine $Q$ on $\mathbb{D}$.
@@
    The representation (1.4c) of $Q$ is not manifestly a polynomial. Rather, as is elaborated in Section 2, it follows from (i) the existence of a solution of Problem 1 via the Féjer-Riesz theorem [18] and (ii) the uniqueness of analytic functions constructed by the Schwarz integral formula that $Q$ is a polynomial.
@@
    Theorem 2 and Corollary 2.2 can be proven in different ways. Consider the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately. One can use (1.1) and the Fejér-Riesz theorem to construct a scalar Riemann-Hilbert problem [19] on $\mathbb{T}$ for $\log(Q(z)/Q_{0}(z))$; this Riemann-Hilbert problem is explicitly solvable using a Cauchy integral, from which (1.4a) follows. Corollary 2.2 can be proven directly using (1.1), the Fejér-Riesz theorem, and the fact that the periodic Hilbert transform (2.11) relates the real and imaginary parts of the boundary values of a function analytic on $\mathbb{D}$, namely $\log(Q(z)/Q_{0}(z))$.
@@ 1.3 Related work
In the quantum signal processing literature, a variety of numerical methods for solving Problem 1 or its avatars have been developed. As is evident from the proof of Theorem 2 in Section 2.1, knowledge of all roots of $1-P(z)P^{*}(1/z)$ allows for the explicit construction of $Q$; analogous statements hold for complementary polynomials in standard QSP. Thus, employing standard root-finding algorithms provides a straightforward means to calculate complementary polynomials [4, 11, 13]. Unfortunately, root-finding algorithms are known to be expensive and suffer from numerical instability, and the highest degree polynomial successfully treated with this approach was reported to have $d=3\times 10^{3}$ [13]. An alternative method that avoids root-finding and instead directly calculates the characteristic polynomial of the roots of $1-P(z)P^{*}(1/z)$ within $\mathbb{D}$ using Prony’s method has been proposed in [15]. Numerical experiments have demonstrated the effectiveness of this approach for polynomials with degree up to $d=5\times 10^{4}$.
The current state-of-the-art method for Problem 1 was developed in [3]. There, a loss function derived from the complementarity condition (1.1) is minimized to determine $Q$ with an optimization procedure; this approach was demonstrated to be effective for $d$ up to the order of $10^{7}$, achieving accuracies as low as $10^{-6}$ in the loss function. In this paper, we present numerical results showing that our algorithm is effective for the same degrees, up to $d=10^{7}$. At the same time our algorithm requires much shorter runtimes and achieves better accuracies, even without the GPU acceleration used in [3].
We mention in passing that as an alternative to the approaches described above, optimization-based methods to compute phase factors without knowledge of the complementary polynomial have been developed in [14, 20, 21]. These methods have been used to determine phase factors for polynomials up to degree $d=10^{5}$.
@@ 1.4 Notation
We write the complex conjugate of $z\in\mathbb{C}$ as $z^{*}$ and for a function $f:\mathbb{C}\to\mathbb{C}$, define $f^{*}(z)\coloneqq f(z^{*})^{*}$. For a set $X\subset\mathbb{C}$, we write $\partial X$ and $\overline{X}$ for its boundary and closure, respectively. We define $\mathbb{D}$ and $\mathbb{T}$ to be open unit disk and unit circle in the complex plane, respectively. Given an integer $N\in\mathbb{Z}_{>0}$, we define the sets $[N]\coloneqq\{n\in\mathbb{Z}_{>0}:n\leq N\}$ and $[N]_{0}\coloneqq[N]\cup\{0\}$. Dashed integrals indicate a principal value prescription with respect to singularities of the integrand on the integration contour. Contour integrals are always assumed to carry a positive orientation. Unless otherwise indicated, all logarithms are with respect to base $\mathrm{e}$. We denote by $\lVert\cdot\rVert_{\infty,\mathbb{T}}$ the uniform norm on $\mathbb{T}$.

### v2

@@ Abstract
Quantum signal processing is a framework for implementing polynomial functions on quantum computers. To implement a given polynomial $P$, one must first construct a corresponding complementary polynomial $Q$. Existing approaches to this problem employ numerical methods that are not amenable to explicit error analysis. We present a new approach to complementary polynomials using complex analysis. Our main mathematical result is a contour integral representation for a canonical complementary polynomial. On the unit circle, this representation has a particularly simple and efficacious Fourier analytic interpretation, which we use to develop a Fast Fourier Transform-based algorithm for the efficient calculation of $Q$ in the monomial basis with explicit error guarantees. Numerical evidence that our algorithm outperforms the state-of-the-art optimization-based method for computing complementary polynomials is provided.
@@ 1 Introduction
Quantum signal processing (QSP) [1] and its extensions [2, 3] describe simple single-qubit parameterized circuits that apply a chosen polynomial to a scalar. They have become indispensable in state-of-the-art quantum algorithms because the circuits can be lifted from scalars to arbitrary matrices, resulting in the quantum singular value transformation (QSVT) [4, 5]. The matrix is embedded inside a larger unitary, called a block encoding. The polynomial is then applied to all singular values of the matrix simultaneously, which can lead to a quantum speedup compared to classical evaluation of matrix functions. The QSVT has had a tremendous impact in quantum computing since polynomials may be used to approximate a wide variety of functions. Thereby, this family of algorithms encompasses many prior quantum algorithms [6], including those for Hamiltonian simulation [7], solving linear systems [8], phase estimation [9], and amplitude amplification [10], often even improving the prior algorithm.
There are different parameterizations of QSP [1, 4, 11]. The fundamental idea underlying each is that certain polynomials may be realized within a matrix element of a finite product of single-qubit unitaries [1, 3, 12]. Determining the parameters of these unitaries is the obstruction to implementing particular polynomials in QSP. To delineate the mathematical structure of this factorization problem, we recall the generalized quantum signal processing (GQSP) framework [^1] For future convenience, in particular to relate (1.2) to other QSP conventions in Appendix A, we have written (1.2) as the transpose of [3, Eq. 7]. of Motlagh and Wiebe [3], which, as we show in Appendix A, subsumes standard formulations of QSP.
@@
In (1.2), we call $Q$ a complementary polynomial to $P$ and the parameters $\lambda$, $(\phi_{j})_{j=0}^{d}$, $(\theta_{j})_{j=0}^{d}$ phase factors. Given a complementary polynomial, the phase factors may be constructed via an exact iterative method [3, Algorithm 1]; analogous statements hold for standard QSP, see [4, Theorem 3] and [13, Section 3.1.2]. Importantly, the phase factors required in lifted algorithms operating on matrices like the QSVT follow immediately from those obtained in QSP. Motivated by the above discussion, we introduce the following problem, whose theoretical and numerical resolution is the subject of this paper.
@@
Previous approaches to this problem rely on root-finding [4], Prony’s method [13], or optimization [3]; see Section 1.3. Here, we construct an exact representation of a canonical complementary polynomial $Q$, valid throughout the entire complex plane, in the form of a set of contour integrals; the problem of constructing the complementary polynomial is thus reduced to quadratures. The contour integral representation of $Q$ on $\mathbb{T}$ can be rephrased in the language of Fourier analysis. We combine this Fourier analytic interpretation and the Fast Fourier Transform (FFT) to develop efficient numerical algorithms to compute $Q$ in the monomial basis.
Exact and explicit error analysis of our algorithms is performed, providing rigorous upper bounds on the classical runtimes. Furthermore, numerical results from our reference implementation demonstrate the practicality and competitiveness of our algorithm. We emphasize that existing numerical approaches to the construction of complementary polynomials, which we describe in Section 1.3 below, rely on heuristics and so are not amenable to rigorous error analysis.
In the remainder of this introduction, we state our results, give remarks on their proofs, describe related literature, introduce notation used in the main text, and outline the structure of the paper.
@@ 1.1 Statement of results
the restriction that $p_{0}\neq 0$ is imposed without loss of generality as $\lvert z^{n}P(z)\rvert=\lvert P(z)\rvert$ holds for all $z\in\mathbb{T}$, $n\in\mathbb{Z}$. Our main result is the following theorem, giving representations of $Q$ on $\mathbb{D}\coloneqq\{z\in\mathbb{C}:\lvert z\rvert<1\}$, $\mathbb{T}$, and $\mathbb{C}\setminus\overline{\mathbb{D}}$.
@@
the integration contour $\mathbb{T}$ is positively-oriented, and the dashed integral indicates a Cauchy principal value prescription (B.4) with respect to the singularity $z^{\prime}=z$ on $\mathbb{T}$, solves Problem 1. Moreover, (1.4c) is, up to a multiplicative phase, the unique solution of Problem 1 with no roots in $\mathbb{D}$.
Theorem 2 provides an exact representation of $Q$ in Problem 1, in canonical form: all roots lie outside of $\mathbb{D}$. In the important special case where $P$ has real coefficients, a canonical complementary polynomial with real coefficients exists. The precise statement, a mild generalization of a result in [14, Section 8], is now given.
@@
Numerical methods.
Corollary 2.2 suggests a practical numerical method to compute an approximation of the complementary polynomial, supposing $Q_{0}$ is known. This is trivially the case if
$\lVert P(z)\rVert_{\infty,\mathbb{T}}\coloneqq\max_{z\in\mathbb{T}}\,\lvert P(z)\rvert\leq 1-\delta\quad(\delta\in(0,1));$ (1.9)
then, $Q_{0}(z)=1$. If $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1$ is guaranteed but a tighter bound (1.9) either does not exist or is unknown, $Q_{0}(z)=1$ can be attained by slightly rescaling $P(z)\to(1-\delta)P(z)$ for a suitable $\delta\in(0,1)$.
In cases with a known bound (1.9), Algorithm 1 solves Problem 1 in time $O(N\log N)$ using a sequence of FFTs, where the even parameter $N\in\mathbb{Z}_{\geq d}$ defines the discrete Fourier basis $\{\mathrm{e}^{\mathrm{i}n\theta}\}_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$. An informal description of our algorithm, based on Corollary 2.2 with $Q_{0}(z)=1$, is as follows.
@@
    Compute approximations to the Fourier coefficients $(a_{n})_{n=-\frac{N}{2}+1}^{\frac{N}{2}}$ of $\log\big{(}1-\lvert P(\mathrm{e}^{\mathrm{i}\theta})\rvert^{2}\big{)}$ using an FFT, in time $O(N\log N)$.
@@
    Compute approximations to $Q$ at the $N$th roots of unity by applying the Fourier multiplier $\Pi$ in Fourier space (1.6)–(1.8), using an FFT and inverse FFT in time $O(N\log N)$.
@@
    Compute approximations to the coefficients of $Q/{Q}_{0}$ in the monomial basis using the result of the previous step and an FFT, in time $O(N\log N)$.
We prove in Theorem 3 that this algorithm is efficient, with a sufficient $N$ scaling as $N=O\!\left(\frac{d}{\delta}\log\frac{d}{\delta\varepsilon}\right)$, where $\varepsilon$ is the error in the in the monomial basis coefficients.
In the case where $\lVert P(z)\rVert_{\infty,\mathbb{T}}\leq 1$, but a tighter upper bound (1.9) either (i) does not exist, (ii) is unknown, or (iii) has $\delta$ so small that the upper bound on the runtime of Algorithm 1 is undesirable, we recourse to Algorithm 2. In this algorithm, the input polynomial is appropriately downscaled and then input into Algorithm 1. Theorem 4 proves that a sufficient $N=O\!\left(\frac{d}{\varepsilon}\log\frac{d}{\varepsilon}\right)$, where $\varepsilon$ is the error in the complementarity condition (1.1); observe that $N$ is independent of $\delta$.
We give a reference implementation of Algorithm 1 and provide numerical evidence that it outperforms the optimization-based approach to complementary polynomials from [3] in Section 5.
@@ 1.2 Remarks on the results and their proofs
The following remarks apply to Theorem 2 and Corollary 2.2. Below, we reference classical complex analysis theorems; precise statements of these theorems may be found in Appendix B.
@@
    The crucial observation leading to Theorem 2 is that the real part of the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately, can be determined exactly on $\mathbb{T}$ using (1.1). Then, the Schwarz integral formula [15] is used to construct $Q$ on $\mathbb{D}$.
@@
    The representation (1.4c) of $Q$ is not manifestly a polynomial. Rather, as elaborated in Section 2, it follows from (i) the existence of a canonical solution of Problem 1 via the Féjer-Riesz theorem [16] and (ii) the uniqueness of holomorphic functions constructed by the Schwarz integral formula that $Q$ is a polynomial.
@@
    Theorem 2 and Corollary 2.2 can be proven in different ways. Consider the function $\log(Q(z)/Q_{0}(z))$, where $Q$ is chosen so that all roots lie outside $\overline{\mathbb{D}}$ and the branch cuts are chosen appropriately. One can use (1.1) and the Fejér-Riesz theorem to construct a scalar Riemann-Hilbert problem [17] on $\mathbb{T}$ for $\log(Q(z)/Q_{0}(z))$; this Riemann-Hilbert problem is explicitly solvable using a Cauchy integral, from which (1.4a) follows. Corollary 2.2 can be proven directly using (1.1), the Fejér-Riesz theorem, and the fact that the periodic Hilbert transform (2.11) relates the real and imaginary parts of the boundary values of a function holomorphic on $\mathbb{D}$, namely $\log(Q(z)/Q_{0}(z))$.
@@ 1.3 Related work
In the QSP literature, a variety of numerical methods for solving Problem 1 or its avatars have been developed. As is evident from the proof of Theorem 2 in Section 2.1, knowledge of all roots of $1-P(z)P^{*}(1/z)$ allows for the explicit construction of $Q$; analogous statements hold for complementary polynomials in standard QSP. Thus, employing standard root-finding algorithms provides a straightforward means to calculate complementary polynomials [4, 11, 14]. Root-finding algorithms are known to be expensive and suffer from numerical instability; the highest-degree polynomial successfully treated with this approach was reported to have degree $d=3\times 10^{3}$ [14]. An alternative method that avoids root-finding and instead directly calculates the characteristic polynomial of the roots of $1-P(z)P^{*}(1/z)$ within $\mathbb{D}$ using Prony’s method has been proposed in [13]. Numerical experiments have demonstrated the effectiveness of this approach for polynomials with degree up to $d=5\times 10^{4}$.
The current state-of-the-art method for Problem 1 was developed in [3]. There, a loss function derived from the complementarity condition (1.1) is minimized to determine $Q$ with an optimization procedure; this approach was demonstrated to be effective for $d$ up to the order of $10^{7}$, achieving accuracies as low as $10^{-6}$ in the loss function. In this paper, we present numerical results showing that Algorithm 1 is effective for the same degrees, up to $d=10^{7}$. At the same time, our algorithm requires much shorter runtimes and achieves better accuracies, even without the GPU acceleration used in [3].
Given a complementary polynomial, it remains to calculate the phase factors. In both GQSP [3] and standard QSP [4, 13], exact recursive formulas may be used to determine phase factors. Variations on and improvements to this approach are described in [11] and [14]. We also mention that, as an alternative to the approaches described above, optimization-based methods to compute phase factors without knowledge of the complementary polynomial have been developed in [18, 19, 20]. These methods have been used to determine phase factors for polynomials up to degree $d=10^{5}$.
@@ 1.4 Notation
We write the complex conjugate of $z\in\mathbb{C}$ as $z^{*}$ and for a function $f:\mathbb{C}\to\mathbb{C}$, define $f^{*}(z)\coloneqq f(z^{*})^{*}$. For a set $X\subset\mathbb{C}$, we write $\partial X$ and $\overline{X}$ for its boundary and closure, respectively. We define $\mathbb{D}$ and $\mathbb{T}$ to be the open unit disk and unit circle in the complex plane, respectively. Given an integer $N\in\mathbb{Z}_{\geq 1}$, we define the sets $[N]\coloneqq\{n\in\mathbb{Z}_{\geq 1}:n\leq N\}$ and $[N]_{0}\coloneqq[N]\cup\{0\}$. Dashed integrals indicate a Cauchy principal value prescription with respect to singularities of the integrand on the integration contour. Contour integrals are always assumed to carry a positive orientation. Unless otherwise indicated, all logarithms are with respect to base $\mathrm{e}$. We denote by $\lVert\cdot\rVert_{\infty,X}$ the uniform norm on $X$.